Please help me answer this question:
Tom's father is 3 times old as Tom. 4 years ago, he was 4 times older. How old is Tom?

Answer by ptaylor(2198)

Let x=Tom's age now
Then 3x=Tom's father's age now
x-4=Tom's age 4 years ago
4(x-4)=Tom's father's age 4 years ago
So, our equation to solve is:
4(x-4)=3x-4 get rid of parens
4x-16=3x-4 subtract 4x from and add 4 to each side
4x-4x-16+4=3x-4x-4+4 collect like terms
-12=-x or
x=12 --------------Tom's age now
3x=3*12=36 Tom's father's age now
CK
4(12-4)=3*12-4
32=36-4
32=32
Hope this helps---ptaylor